An elephant lifts a fallen tree by exerting a force of 8.35 x 103 N upward. If she lifts the tree 1.95 m in 1.25 seconds, how much power does she use?



1
Expert's answer
2020-12-30T20:41:43-0500

Power is how quickly work can be done:


"P=\frac Wt."

The work in our case was spent on changing potential energy of the tree from zero to the final value:


"W=mgh=Fh."

Therefore:


"P=\frac{Fh}{t}=\frac{8.35\cdot10^3\cdot1.95}{1.25}=13062\text{ W}."